# Strong coalitions in graphs

### v1

@@ Abstract
Keywords: Coalition; strong coalition, bound.
@@ 1 Introduction
Domination in graphs is well studied in graph theory and the literature on this subject has been surveyed and detailed in the books [12, 13]. Given a graph $G=(V,E)$, recall that a dominating set $S$ of a graph $G$ is a set such that every vertex in $V\setminus S$ is adjacent to some vertex in $S$. The domination number $\gamma(G)$ is the minimum cardinality of a dominating set. A set $D\subset V$ is a strong dominating set of $G$, if for every vertex $x\in V\setminus D$ there is a vertex $y\in D$ with $xy\in E(G)$ and $deg(x)\leq deg(y)$. The strong domination number $\gamma_{st}(G)$ is defined as the minimum cardinality of a strong dominating set. The strong domination number was introduced in [15] and has been studied in [4, 14, 16]. A domatic partition (or strong partition) is a partition of the vertex set into dominating sets (or strong dominating sets). The maximum cardinality of a domatic partition (or strong partition) is called the domatic number (or strong domatic number), denoted by $d(G)$ (or $d_{st}(G)$). The domatic number of a graph was introduced in 1977 by Cockayne and Hedetniemi [5]. Very recently, the strong domatic of a graph has started to be explored in [6]. For more details on the domatic number refer to e.g., [17, 18].
A coalition in a graph $G=(V,E)$ consists of two disjoint sets $V_{1}$ and $V_{2}$ of vertices, such that neither $V_{1}$ nor $V_{2}$ is a dominating set, but the union $V_{1}\cup V_{2}$ is a dominating set of $G$. A coalition partition in a graph $G$ of order $n=|V|$ is a vertex partition $\mathcal{P}=\{V_{1},V_{2},...,V_{k}\}$ such that every set $V_{i}$ either is a dominating set consisting of a single vertex of degree $n-1$, or is not a dominating set but forms a coalition with another set $V_{j}$ which is not a dominating set. In 2020, Hedetniemi et al. [7] introduced and studied the concept of coalitions in graphs. For some recent papers on some variants of coalitions in graphs see [1, 2, 3, 8, 9, 10, 11]. In order to expand the study of coalitions, we consider coalitions involving strong dominating sets in graphs.
@@ 2 Existence of strong coalition partitions
A strong coalition partition of a graph $G=(V,E)$ is a partition $\Omega=\{V_{1},V_{2},\ldots,V_{k}\}$ of the vertex set $V$ such that any $V_{i}\in\Omega,1\leq i\leq k,$ is either a strong dominating set comprising a single vertex of degree $n-1$, or is not a strong dominating set but forms a strong coalition with another set $V_{j}\in\Omega$ that is not a strong dominating set. The maximum number of sets in $\Omega$ gives the strong coalition number $SC(G)$ of $G$. A partition of the vertex set $V$ into $SC(G)$ strong sets is called a $SC(G)$-partition.
The main goal of this section is to find out a necessary condition for an existence of a strong coalition partition in a graph $G$. To achieve this goal, we shall now construct families of graphs have no strong coalition partitions.
Observation 2.3
  1. (i)
    The star graphs $K_{1,n}$, for $n\geq 2$ have no strong coalition partitions.
  2. (ii)
    The friendship graphs $F_{n}$ have no strong coalition partitions.
Let the family of graphs ${\cal F}$ contains all graphs $G\neq K_{n}$ of order $n$ with at least one vertex of degree $n-1$. The next result provides a necessary condition for the existence of a strong coalition partition.
@@ 3 Bounds on the strong coalition number
If $G$ is a graph of order $n$ with no full vertex, then $SC(G)\geq 2d_{st}(G)$.
@@
If $G$ is a graph of order $n$ with maximum degree $\Delta$, then
$SC(G)\leq 2+2\Delta.$
Proof.  If $\Delta(G)\geq\frac{n-1}{2}$, then by Corollary 3.5 we have the result. Now, suppose that $\Delta(G)<\frac{n-1}{2}$. Let $V(G)=\{v_{1},v_{2},\ldots,v_{n}\}$ and $\Omega=\{V_{1},V_{2},\ldots,V_{t}\}$ be a $SC(G)$-partition. Every set of $\Omega$ that does not contain either $v_{1}$ and $v_{2}$ must be in a strong coalition with a set containing either $v_{1}$ and $v_{2}$. If $v_{1}$ and $v_{2}$ are in the same set of $\Omega$, then Theorem 3.6 implies that $SC(G)\leq 1+1+\Delta(G)$. So, clearly $SC(G)\leq 2+2\Delta(G)$. Now, suppose that $v_{1}$ and $v_{2}$ are in different sets of $\Omega$. Without loss of generality, let $v_{1}\in V_{1}$ and $v_{2}\in V_{2}$. So, each of $V_{1}$ and $V_{2}$ is in a strong coalition with at most $1+\Delta(G)$ sets. If $V_{1}$ and $V_{2}$ form a strong coalition, then $SC(G)\leq 2+2\Delta(G)$ as each of $V_{1}$ and $V_{2}$ can be in at most $\Delta(G)$ additional strong coalitions. Otherwise, if $V_{1}$ and $V_{2}$ does not form a strong coalition, we also have the result. To show it, suppose that $t>2+2\Delta(G)$. Without loss of generality, we may assume that $V_{1}$ is in a strong coalition with each of $V_{3}$, $V_{4}$, $\ldots$, $V_{\Delta(G)+2}$ and $V_{\Delta(G)+3}$, while $V_{2}$ is in a strong coalition with each of $V_{\Delta(G)+4}$, $V_{\Delta(G)+5}$, $\ldots$, $V_{2\Delta(G)+2}$ and $V_{2\Delta(G)+3}$. Let $v_{i}$ be a vertex that is not strong dominated by $V_{1}$ and $\deg(v_{i})\leq\Delta(G)$. without loss of generality, suppose that $N(v_{i})=\{v_{1}^{\prime},v_{2}^{\prime},\ldots,v_{\Delta(G)}^{\prime}\}$, $v_{i}\in V_{3}$ , and $v_{1}^{\prime}\in V_{4}$, $v_{2}^{\prime}\in V_{5}$, $\ldots$, $v_{\Delta(G)-1}^{\prime}\in V_{\Delta(G)+2}$ and $v_{\Delta(G)}^{\prime}\in V_{\Delta(G)+3}$. But now $V_{2}\cup V_{\Delta(G)+4}$ does not strong dominate $v_{i}$, contradicting that $V_{2}$ and $V_{\Delta(G)+4}$ form a strong coalition in $G$. $\square$
@@
Proof.  Suppose that $u$ is the vertex with maximum degree. Since $u$ should be in every strong dominating sets like $S$, so $S\setminus\{u\}$ and $\{u\}$ form strong coalition and we are done. $\square$
@@
Study graphs $G$ with large and small strong coalition numbers are interesting problems. In this short section, we give a sufficient condition for graphs $G$ of order at least 4 with $SC(G)=|V(G)|$. To construct a family ${\cal G}$ of graphs with this properties, consider a complete graph $K_{r+s}$, where its vertex set is $A\cup B$ such that $A=\{u_{1},...,u_{r}\}$, $B=\{v_{1},...,v_{s}\}$, $A\cap B=\emptyset$, $r\geq 0$, $s\geq 0$. Any graph $G$ of order at least 4 in the family ${\cal G}$ have (see Figure 1), $V(G)=A\cup B\cup 2V(K_{p})$ ($p\geq 0$), where $V(K_{p})_{1}=\{x_{1},x_{2},...,x_{p}\}$ and $V(K_{p})_{2}=\{w_{1},w_{2},...,w_{p}\}$ and
@@ 5 Strong coalition number for some classes of graphs
By Proposition 5.1 and Observation 5.2, we observe that there are graphs $G$ such that the difference between $C(G)$ and $SC(G)$ can be large enough.
@@
There exist a graph $G$ such that $|C(G)-SC(G)|$ is very large.
@@
For the path $P_{n}$ of order $n\geq 4$,
$SC(P_{n})=\left\{\begin{array}[]{ll}4&\textrm{~{}~{}~{}~{} if ~{}~{}$4\leq n\leq 7$}\\ \\ 5&\textrm{~{}~{}~{}~{} if ~{}~{} $8\leq n\leq 11$}\\ \\ 6&\textrm{~{}~{}~{}~{} if ~{}~{}$n\geq 12$}\end{array}\right.$
@@
Proof.  Let $V(P_{n})=\{v_{1},v_{2},\ldots,v_{n}\}$ such that $\deg(v_{1})=\deg(v_{n})=1$ and $\deg(v_{i})=2$ for $12\leq i\leq n-$ 1 and $v_{i-1}$ and $v_{i}$ are adjacent, for $2\leq i\leq n$. First suppose that $n=4$. Since no single vertex strongly dominates $P_{4}$ but any pair of non-adjacent vertices strongly dominate $P_{4}$, the partition containing each vertex in a singleton set is a $SC(G)$-partition of $P_{4}$ of order $4$. Now assume that $n=5$ and let $\Omega=\{\{v_{1}\},\{v_{2}\},\{v_{3}\},\{v_{4}\},\{v_{5}\}\}$ be a $SC(G)$-partition of $P_{5}$ into singleton sets. Since the graph $P_{5}$ has a unique strong dominating set of order two, i.e., $\{v_{2},v_{4}\}$, then each set containing vertices $v_{1}$, $v_{3}$, and $v_{5}$ is not in a strong coalition with any other set of $\Omega$. So only the sets that contain $v_{2}$ and $v_{4}$ are classified as singleton sets. Thus, in order for $\Omega$ to be a $SC(G)$-partition of $P_{5}$, $\Omega$ must contain at least one non-singleton set. So, $SC(P_{5})\leq 4$. The partition $\{V_{1}=\{v_{3},v_{5}\},V_{2}=\{v_{1}\},V_{3}=\{v_{2}\},V_{4}=\{v_{4}\}\}$ is a $SC(G)$-partition of $P_{5}$, in which $V_{1}$ forms a strong coalition with each of $V_{2}$ and $V_{3}$, and $V_{4}$ forms a strong coalition with $V_{3}$. Now, we consider $n=6$. Let $\Omega=\{\{v_{1}\},\{v_{2}\},\{v_{3}\},\{v_{4}\},\{v_{5}\},\{v_{6}\}\}$ be a $SC(G)$-partition of $P_{6}$. Similar to $P_{5}$, any strong coalition partition of $P_{6}$ must contain at least one non-singleton set. Therefore, $SC(P_{6})\leq 5$. Now, we show that $SC(P_{6})\neq 5$. Let $\Omega=\{V_{1},V_{2},V_{3},V_{4},V_{5}\}$ be a $SC(G)$-partition of $P_{6}$. Suppose that $V_{1}=\{v_{1}\}$. It is clear that $V_{1}$ can not form a strong coalition with a singleton set. It follows that a strong coalition partner of $V_{1}$ must have at least two elements. One can easily check that the only option is $V_{2}=\{v_{3},v_{5}\}$. If other sets be singleton sets, none of them can form a strong coalition with the singleton set consisting of $v_{4}$. The same argument holds for the singleton set consisting of $v_{6}$. So we need at least two sets with two vertices or one set with at least 3 vertices. Therefore in this case $SC(P_{6})\neq 5$. Now assume that $|V_{1}|=2$ and $|V_{i}|=1$, for $2\leq i\leq 5$. Let $V_{1}=\{v_{1},v_{j}\}$ such that $j\in\{2,3,4,5,6\}$. If $i\neq 6$, then the singleton set consisting of $v_{6}$ can only form a strong coalition with the doubleton set consisting of $\{v_{2},v_{4}\}$. If $V_{1}=\{v_{1},v_{6}\}$, then $V_{1}$ must be in a strong coalition with a set with size at least two. So in all cases, there is no $SC(G)$-partition of $P_{6}$ of order 5. One can easily check that the partition $\{V_{1}=\{v_{1},v_{6}\},V_{2}=\{v_{3},v_{4}\},V_{3}=\{v_{2}\},V_{4}=\{v_{5}\}\}$ is a $SC(G)$-partition of $P_{6}$. Next assume that $n=7$. By Theorem 3.7, we have $SC(P_{7})\leq 6$. We show that $SC(P_{7})\neq 6$. Let $\Omega$ be a $SC(G)$-partition of $P_{7}$. Note that $\Omega$ consists of a set, say $V_{1}$, of cardinality $2$ and five singleton sets. Since $\gamma_{st}(P_{7})=3$, each singleton set must form a strong coalition with $V_{1}$, and by Theorem 3.6, $V_{1}$ has at most three strong coalition partners, a contradiction. Therefore, $SC(P_{7})\leq 5$. With similar method we see that $SC(P_{7})\neq 5$. The partition $\{V_{1}=\{v_{1},v_{3}\},V_{2}=\{v_{2},v_{4}\},V_{3}=\{v_{6}\},V_{4}=\{v_{5},v_{7}\}\}$ is a $SC(G)$-partition of $P_{7}$. Next suppose that $n=8$. By Theorem 3.7, we know that $SC(P_{8})\leq 6$. We first show that $SC(P_{8})\neq 6$. Let $\Omega$ be a $SC(G)$-partition of $P_{8}$. Now, we consider two cases.
@@
Case 1. $|V_{1}|=5$ and $|V_{i}|=1$ for $2\leq i\leq 6$. Since $\gamma_{st}(P_{10})=4$, no two singleton sets in $\Omega$ can produce strong coalitions. Therefore, each singleton set must produce a strong coalition with $V_{1}$. By Theorem 3.6, $V_{1}$ is in at most three strong coalition partners, a contradiction. So, $SC(P_{10})\leq 5$.
Case 2. $|V_{1}|=4$, $|V_{2}|=2$, and $|V_{i}|=1$ for $3\leq i\leq 6$. Since $\gamma_{st}(P_{10})=4$, no two singleton sets in $\Omega$ can form strong coalitions. In addition, no singleton set forms a strong coalition with $V_{2}$. Thus, each of $V_{j}$ for $2\leq j\leq 6$ must be in a strong coalition with $V_{1}$. By Theorem 3.6, $V_{1}$ has at most three strong coalition partners, a contradiction. Hence, $SC(P_{10})\leq 5$.
@@
Subcase (i). Let $V_{2}$ and $V_{3}$ do not form a strong coalition. Therefore, each of $V_{j}$ for $2\leq j\leq 6$ must form a strong coalition with $V_{1}$, a contradiction. Hence, $SC(P_{10})\leq 5$.
@@
Case 1.$|V_{1}|=6$ and $|V_{i}|=1$ for $2\leq i\leq 6$. Since $\gamma_{st}(P_{11})=4$, no two singleton sets in $\Omega$ can form strong coalitions. Thus, each singleton set must produce a strong coalition with $V_{1}$. By Theorem 3.6, $V_{1}$ has at most three strong coalition partners, a contradiction. Hence, $SC(P_{11})\leq 5$.
@@
Case 2.$|V_{1}|=5$, $|V_{2}|=2$, and $|V_{i}|=1$ for $3\leq i\leq 6$. Since $\gamma_{st}(P_{11})=4$, each of $|V_{j}|=1$ for $2\leq j\leq 6$ must be in a strong coalition with $V_{1}$. By Theorem 3.6, $V_{1}$ is in at most three strong coalition partners, a contradiction. So, $SC(P_{11})\leq 5$.
@@
Case 3. $|V_{1}|=4,|V_{2}|=3$, and $|V_{i}|=1$ for $3\leq i\leq 6$. Since $\gamma_{st}(P_{11})=4$, each singleton set must produce a strong coalition with $V_{1}$ or $V_{2}$. It can be seen that $P_{11}$ has only two strong coalitions of order 4 such as $A=\{v_{2},v_{5},v_{8},v_{10}\}$ and $B=\{v_{2},v_{5},v_{7},v_{10}\}$. Now, we consider the following subcases.
@@
Subcase (i). Without loss of generality, assume that $V_{2}\cup V_{3}=\{v_{2},v_{5},v_{7},v_{10}\}$ is a strong coalition of order 4. As be seen, each of $V_{4}$, $V_{5}$, and $V_{6}$ must be in a strong coalition with $V_{1}$, which is impossible, as the pairs $V_{1}\cup V_{4}$, $V_{1}\cup V_{5}$, and $V_{1}\cup V_{6}$ that comprise the remaining vertices $v_{1},v_{3},v_{4},v_{5},v_{8}$, and $v_{9}$ have no three vertices in common. So, $SC(P_{11})\leq 5$.
@@
Subcase (ii). Let $V_{2}\cup V_{3}=\{v_{2},v_{5},v_{8},v_{10}\}$. As be seen, each of $V_{4}$, $V_{5}$, and $V_{6}$ must be in a strong coalition with $V_{1}$, which is impossible, as the pairs $V_{1}\cup V_{4}$, $V_{1}\cup V_{5}$, and $V_{1}\cup V_{6}$ that comprise the remaining vertices $v_{1},v_{3},v_{4},v_{5},v_{7}$, and $v_{9}$ have no three vertices in common. So, $SC(P_{11})\leq 5$.

### v2

@@ Abstract
Keywords: Coalition, strong dominating set, strong coalition.
@@ 1 Introduction
We begin with some basic definitions. Let $G$ be a graph with vertex set $V=V(G)$, edge set $E=E(G)$, and order $n=|V|$. The minimum and maximum degrees in $G$ are denoted by $\delta(G)$ and $\Delta(G)$, respectively. A universal vertex in a graph $G$ is a vertex adjacent to all other vertices, and so $deg(v)=n-1$. Domination is an extensively studied classic topic in graph theory and the literature on this subject has been thoroughly reviewed in the books [14, 15]. Given a graph $G=(V,E)$, recall that a dominating set $S$ of a graph $G$ is a set such that every vertex in $V\setminus S$ is adjacent to some vertex in $S$. The domination number $\gamma(G)$ is the minimum cardinality of a dominating set. A set $D\subset V$ is a strong dominating set of $G$, if for every vertex $x\in V\setminus D$ there is a vertex $y\in D$ with $xy\in E(G)$ and $deg(x)\leq deg(y)$. The strong domination number $\gamma_{st}(G)$ is defined as the minimum cardinality of a strong dominating set. The strong domination number was introduced in [17] inspired by the real-life application to network traffic, and has subsequently been studied in [1, 5, 8, 16, 18]. A domatic partition (or strong partition) is a partition of the vertex set into dominating sets (or strong dominating sets). The maximum cardinality of a domatic partition (or strong partition) is called the domatic number (or strong domatic number), denoted by $d(G)$ (or $d_{st}(G)$). The domatic number of a graph was introduced in 1977 by Cockayne and Hedetniemi [7]. Very recently, the strong domatic number of a graph has started to be explored in [9]. For more details on the domatic number refer to e.g., [19, 20].
A coalition in a graph $G=(V,E)$ consists of two disjoint sets $V_{1}$ and $V_{2}$ of vertices, such that neither $V_{1}$ nor $V_{2}$ is a dominating set, but the union $V_{1}\cup V_{2}$ is a dominating set of $G$. A coalition partition in a graph $G$ of order $n=|V|$ is a vertex partition $\mathcal{P}=\{V_{1},V_{2},...,V_{k}\}$ such that every set $V_{i}$ either is a dominating set consisting of a single vertex of degree $n-1$, or is not a dominating set but forms a coalition with another set $V_{j}$ which is not a dominating set. The maximum cardinality of a coalition partition is the coalition number of a graph $G$, and is denoted by $C(G)$. In 2020, Hedetniemi et al. [10] introduced and studied the concept of coalition in graphs and have been investigated in [6, 11, 12, 13]. For some recent papers on some variants of coalitions in graphs see [2, 3, 4]. In order to expand the study of coalitions, we consider coalitions involving strong dominating sets in graphs.
@@ 2 Existence of strong coalition partitions
A strong coalition partition of a graph $G=(V,E)$ is a partition $\Omega=\{V_{1},V_{2},\ldots,V_{k}\}$ of the vertex set $V$ such that any $V_{i}\in\Omega,1\leq i\leq k,$ is either a strong dominating set comprising a single vertex of degree $n-1$, or is not a strong dominating set but forms a strong coalition with another set $V_{j}\in\Omega$ that is not a strong dominating set. The maximum number of sets in $\Omega$ gives the strong coalition number $SC(G)$ of $G$.
The main goal of this section is to find out a necessary condition for the existence of a strong coalition partition in a graph $G$. To achieve this goal, we shall now construct families of graphs have no strong coalition partitions. Let the family of graphs ${\cal F}$ contains all graphs $G\neq K_{n}$ of order $n$ with at least one vertex of degree $n-1$. The next result provides a necessary condition for the existence of a strong coalition partition.
@@
As an immediate consequence of Proposition 2.3, we have the following observation.
Observation 2.4
  1. (i)
    If $G$ is a star graph $K_{1,n}$ where $n\geq 2$, then $SC(G)=0$.
  2. (ii)
    If $F_{n}$ is a friendship graph, then $SC(G)=0$.
@@ 3 Bounds on the strong coalition number
If $G$ is a graph of order $n$ contains no universal vertex, then $SC(G)\geq 2d_{st}(G)$.
@@
For any graph $G$ with $\delta(G)=1$, we have $SC(G)\leq 2+2\Delta(G).$
Proof.  If $\Delta(G)\geq\frac{n-1}{2}$, then by Corollary 3.5 we have the result. Now, suppose that $\Delta(G)<\frac{n-1}{2}$. Let $V(G)=\{v_{1},v_{2},\ldots,v_{n}\}$ and $\Omega=\{V_{1},V_{2},\ldots,V_{t}\}$ be a $SC(G)$-partition. Suppose that $\deg(v_{1})=1$. Every set of $\Omega$ that does not contain either $v_{1}$ and $v_{2}$ must be in a strong coalition with a set containing either $v_{1}$ and $v_{2}$. If $v_{1}$ and $v_{2}$ are in the same set of $\Omega$, then since $\deg(v_{1})=1$, Theorem 3.6 implies that $SC(G)\leq 1+1+\Delta(G)$. So, clearly $SC(G)\leq 2+2\Delta(G)$. Now, suppose that $v_{1}$ and $v_{2}$ are in different sets of $\Omega$. Without loss of generality, let $v_{1}\in V_{1}$ and $v_{2}\in V_{2}$. So, each of $V_{1}$ and $V_{2}$ is in a strong coalition with at most $1+\Delta(G)$ sets. If $V_{1}$ and $V_{2}$ form a strong coalition, then $SC(G)\leq 2+2\Delta(G)$ as each of $V_{1}$ and $V_{2}$ can be in at most $\Delta(G)$ additional strong coalitions. Otherwise, if $V_{1}$ and $V_{2}$ do not form a strong coalition, we also have the result. To show it, suppose that $t>2+2\Delta(G)$. Without loss of generality, we may assume that $V_{1}$ is in a strong coalition with each of $V_{3}$, $V_{4}$, $\ldots$, $V_{\Delta(G)+2}$ and $V_{\Delta(G)+3}$, while $V_{2}$ is in a strong coalition with each of $V_{\Delta(G)+4}$, $V_{\Delta(G)+5}$, $\ldots$, $V_{2\Delta(G)+2}$ and $V_{2\Delta(G)+3}$. Let $v_{i}$ be a vertex that is not strongly dominated by $V_{1}$ and $\deg(v_{i})\leq\Delta(G)$. Without loss of generality, suppose that $N(v_{i})=\{v_{1}^{\prime},v_{2}^{\prime},\ldots,v_{\Delta(G)}^{\prime}\}$, $v_{i}\in V_{3}$, and $v_{1}^{\prime}\in V_{4}$, $v_{2}^{\prime}\in V_{5}$, $\ldots$, $v_{\Delta(G)-1}^{\prime}\in V_{\Delta(G)+2}$ and $v_{\Delta(G)}^{\prime}\in V_{\Delta(G)+3}$. But now $V_{2}\cup V_{\Delta(G)+4}$ does not strong dominate $v_{i}$, contradicting that $V_{2}$ and $V_{\Delta(G)+4}$ form a strong coalition in $G$. $\square$
@@
Proof.  Suppose that $u$ is the vertex with maximum degree. Since $u$ should be in every strong dominating sets like $S$, so $S\setminus\{u\}$ and $\{u\}$ form a strong coalition and we are done. $\square$
@@
Studying graphs $G$ with large and small strong coalition numbers form interesting problems. In this short section, we give a sufficient condition for graphs $G$ of order at least 4 with $SC(G)=|V(G)|$. To construct a family ${\cal G}$ of graphs with these properties, consider a complete graph $K_{r+s}$, where its vertex set is $A\cup B$ such that $A=\{u_{1},...,u_{r}\}$, $B=\{v_{1},...,v_{s}\}$, $A\cap B=\emptyset$, $r\geq 0$, $s\geq 0$. Any graph $G$ of order at least 4 in the family ${\cal G}$ has (see Figure 1), $V(G)=A\cup B\cup 2V(K_{p})$ ($p\geq 0$), where $V(K_{p})_{1}=\{x_{1},x_{2},...,x_{p}\}$ and $V(K_{p})_{2}=\{w_{1},w_{2},...,w_{p}\}$ and
@@ 5 Strong coalition number for some classes of graphs
By Proposition 5.1 and Observation 5.2, we have the following corollary.
@@
For every integer $m$ there exists a graph $G$ such that $C(G)-SC(G)=m$.
@@
For the path $P_{n}$,
$SC(P_{n})=\left\{\begin{array}[]{ll}1&\textrm{~{}~{}~{}~{} if ~{}~{}$n=1$}\\ \\ 2&\textrm{~{}~{}~{}~{} if ~{}~{}$n=2$}\\ \\ 0&\textrm{~{}~{}~{}~{} if ~{}~{}$n=3$}\\ \\ 4&\textrm{~{}~{}~{}~{} if ~{}~{}$4\leq n\leq 7$}\\ \\ 5&\textrm{~{}~{}~{}~{} if ~{}~{} $8\leq n\leq 11$}\\ \\ 6&\textrm{~{}~{}~{}~{} if ~{}~{}$n\geq 12$}\end{array}\right.$
@@
Proof.  Let $V(P_{n})=\{v_{1},v_{2},\ldots,v_{n}\}$ such that $\deg(v_{1})=\deg(v_{n})=1$ and $\deg(v_{i})=2$ for $12\leq i\leq n-$ 1 and $v_{i-1}$ and $v_{i}$ are adjacent, for $2\leq i\leq n$. By Lemma 3.4, we have $SC(P_{1})=1$. If $n=2$, then Proposition 5.1 gives the result. By Proposition 2.3, we have $SC(P_{3})=0$. Now suppose that $n=4$. Since no single vertex strongly dominates $P_{4}$ but any pair of non-adjacent vertices strongly dominate $P_{4}$, the partition containing each vertex in a singleton set is a $SC(G)$-partition of $P_{4}$ of order $4$. Now assume that $n=5$ and let $\Omega=\{\{v_{1}\},\{v_{2}\},\{v_{3}\},\{v_{4}\},\{v_{5}\}\}$ be a $SC(G)$-partition of $P_{5}$ into singleton sets. Since the graph $P_{5}$ has a unique strong dominating set of order two, i.e., $\{v_{2},v_{4}\}$, then each set containing vertices $v_{1}$, $v_{3}$, and $v_{5}$ is not in a strong coalition with any other set of $\Omega$. So only the sets that contain $v_{2}$ and $v_{4}$ are classified as singleton sets. Thus, in order for $\Omega$ to be a $SC(G)$-partition of $P_{5}$, $\Omega$ must contain at least one non-singleton set. So, $SC(P_{5})\leq 4$. The partition $\{V_{1}=\{v_{3},v_{5}\},V_{2}=\{v_{1}\},V_{3}=\{v_{2}\},V_{4}=\{v_{4}\}\}$ is a $SC(G)$-partition of $P_{5}$, in which $V_{1}$ forms a strong coalition with each of $V_{2}$ and $V_{3}$, and $V_{4}$ forms a strong coalition with $V_{3}$. Now, we consider $n=6$. Let $\Omega=\{\{v_{1}\},\{v_{2}\},\{v_{3}\},\{v_{4}\},\{v_{5}\},\{v_{6}\}\}$ be a $SC(G)$-partition of $P_{6}$. Similar to $P_{5}$, any strong coalition partition of $P_{6}$ must contain at least one non-singleton set. Therefore, $SC(P_{6})\leq 5$. Now, we show that $SC(P_{6})\neq 5$. Let $\Omega=\{V_{1},V_{2},V_{3},V_{4},V_{5}\}$ be a $SC(G)$-partition of $P_{6}$. Suppose that $V_{1}=\{v_{1}\}$. It is clear that $V_{1}$ can not form a strong coalition with a singleton set. It follows that a strong coalition partner of $V_{1}$ must have at least two elements. One can easily check that the only option is $V_{2}=\{v_{3},v_{5}\}$. If other sets be singleton sets, none of them can form a strong coalition with the singleton set consisting of $v_{4}$. The same argument holds for the singleton set consisting of $v_{6}$. So we need at least two sets with two vertices or one set with at least 3 vertices. Therefore in this case $SC(P_{6})\neq 5$. Now assume that $|V_{1}|=2$ and $|V_{i}|=1$, for $2\leq i\leq 5$. Let $V_{1}=\{v_{1},v_{j}\}$ such that $j\in\{2,3,4,5,6\}$. If $i\neq 6$, then the singleton set consisting of $v_{6}$ can only form a strong coalition with the doubleton set consisting of $\{v_{2},v_{4}\}$. If $V_{1}=\{v_{1},v_{6}\}$, then $V_{1}$ must be in a strong coalition with a set with size at least two. So in all cases, there is no $SC(G)$-partition of $P_{6}$ of order 5. One can easily check that the partition $\{V_{1}=\{v_{1},v_{6}\},V_{2}=\{v_{3},v_{4}\},V_{3}=\{v_{2}\},V_{4}=\{v_{5}\}\}$ is a $SC(G)$-partition of $P_{6}$. Next assume that $n=7$. By Theorem 3.7, we have $SC(P_{7})\leq 6$. We show that $SC(P_{7})\neq 6$. Let $\Omega$ be a $SC(G)$-partition of $P_{7}$. Note that $\Omega$ consists of a set, say $V_{1}$, of cardinality $2$ and five singleton sets. Since $\gamma_{st}(P_{7})=3$, each singleton set must form a strong coalition with $V_{1}$, and by Theorem 3.6, $V_{1}$ has at most three strong coalition partners, a contradiction. Therefore, $SC(P_{7})\leq 5$. With similar method we see that $SC(P_{7})\neq 5$. The partition $\{V_{1}=\{v_{1},v_{3}\},V_{2}=\{v_{2},v_{4}\},V_{3}=\{v_{6}\},V_{4}=\{v_{5},v_{7}\}\}$ is a $SC(G)$-partition of $P_{7}$. Next suppose that $n=8$. By Theorem 3.7, we know that $SC(P_{8})\leq 6$. We first show that $SC(P_{8})\neq 6$. Let $\Omega$ be a $SC(G)$-partition of $P_{8}$. Now, we consider two cases.
@@
Case 1. $|V_{1}|=5$ and $|V_{i}|=1$ for $2\leq i\leq 6$. Since $\gamma_{st}(P_{10})=4$, no two singleton sets in $\Omega$ can produce strong coalitions. Then each singleton set must produce a strong coalition with $V_{1}$, contradicting Theorem 3.6. So, $SC(P_{10})\leq 5$.
Case 2. $|V_{1}|=4$, $|V_{2}|=2$, and $|V_{i}|=1$ for $3\leq i\leq 6$. Since $\gamma_{st}(P_{10})=4$, no two singleton sets in $\Omega$ can form strong coalitions. In addition, no singleton set forms a strong coalition with $V_{2}$. Therefore each of $V_{j}$ for $2\leq j\leq 6$ must be in a strong coalition with $V_{1}$, contradicting Theorem 3.6. Hence, $SC(P_{10})\leq 5$.
@@
Subcase (i). Let $V_{2}$ and $V_{3}$ do not form a strong coalition. Thus each of $V_{j}$ for $2\leq j\leq 6$ must form a strong coalition with $V_{1}$, contradicting Theorem 3.6. Hence, $SC(P_{10})\leq 5$.
@@
Case 1.$|V_{1}|=6$ and $|V_{i}|=1$ for $2\leq i\leq 6$. Since $\gamma_{st}(P_{11})=4$, no two singleton sets in $\Omega$ can form strong coalitions. Then each singleton set must produce a strong coalition with $V_{1}$, which contradicts Theorem 3.6. Hence, $SC(P_{11})\leq 5$.
@@
Case 2.$|V_{1}|=5$, $|V_{2}|=2$, and $|V_{i}|=1$ for $3\leq i\leq 6$. Since $\gamma_{st}(P_{11})=4$, each of $|V_{j}|=1$ for $2\leq j\leq 6$ must be in a strong coalition with $V_{1}$, and by Theorem 3.6, $V_{1}$ is in at most three strong coalition partners, a contradiction. So, $SC(P_{11})\leq 5$.
@@
Case 3. $|V_{1}|=4,|V_{2}|=3$, and $|V_{i}|=1$ for $3\leq i\leq 6$. Since $\gamma_{st}(P_{11})=4$, each singleton set must produce a strong coalition with $V_{1}$ or $V_{2}$. It can be seen that $P_{11}$ possesses only three strong coalitions of order 4 such as $A=\{v_{2},v_{5},v_{8},v_{10}\}$, $B=\{v_{2},v_{5},v_{7},v_{10}\}$, and $C=\{v_{2},v_{4},v_{7},v_{10}\}$. Now, we consider the following subcases.
@@
Subcase (i). Without loss of generality, assume that $V_{2}\cup V_{3}=\{v_{2},v_{5},v_{7},v_{10}\}$ is a strong coalition of order 4. Note that each of $V_{4}$, $V_{5}$, and $V_{6}$ is in a strong coalition with $V_{1}$. It holds that $\bigcap_{k=4}^{6}|V_{1}\cup V_{k}|=4$, which is impossible, as the pairs $V_{1}\cup V_{4}$, $V_{1}\cup V_{5}$, and $V_{1}\cup V_{6}$ comprising the remaining vertices $v_{1},v_{3},v_{4},v_{6},v_{8}$, $v_{9}$ and $v_{11}$ have no four vertices in common. So, $SC(P_{11})\leq 5$.
@@
Subcase (ii). Let $V_{2}\cup V_{3}=\{v_{2},v_{5},v_{8},v_{10}\}$. Assume that each of $V_{4}$, $V_{5}$, $V_{6}$ can be a strong coalition with $V_{1}$. Then $\bigcap_{k=4}^{6}|V_{1}\cup V_{k}|=4$, which is impossible, as the pairs $V_{1}\cup V_{4}$, $V_{1}\cup V_{5}$, and $V_{1}\cup V_{6}$ containing the remaining vertices have no four vertices in common. Hence, $SC(P_{11})\leq 5$.